How many molecules are in 127.39 g of SO3?


1
Expert's answer
2021-02-01T03:55:44-0500

n(SO3)=127.39/80.066=1.591 mol

N(SO3)= 1.591*6.022*1023=9.581*1023 molecules

Answer: 9.581*1023 molecules